OpenStudy (hugsnotughs):

I started to solve a radical equation in this way...but both solutions are extraneous because they don't satisfy the original equation...what error did I make?

OpenStudy (hugsnotughs):

http://prntscr.com/4ll41e

OpenStudy (bibby):

You can't do something like \(\huge \sqrt{-2x+1}=x+3\) \(\huge \sqrt{-2x+1-1}=x+3-1\) \(\huge \sqrt{-2x}=x+2\)

OpenStudy (bibby):

square both sides to cancel the root

OpenStudy (hugsnotughs):

So my mistake was subtracting the ones before squaring the root?

OpenStudy (hugsnotughs):

Thank you.

OpenStudy (bibby):

correct, you'd have to do something like \(\huge \sqrt{-2x+1}^2={(x+3)}^2\)

OpenStudy (hugsnotughs):

Oh, okay. Thanks! You're much help! ;D